A weak acid (HA) has a pKa of 4.679. If a solution of this acid has a pH of 4.876, what percentage of the acid is not ionized? (Assume all H in solution came from the ionization of HA.)
Answer #1
Concepts and reason

Problem is based on the concept of chemical equilibrium. Equilibrium is attained by a reaction if rate of change of concentration of reactant is equal to rate of change of concentration of product.

For a solution, pH is the measure of acidity/basicity of a compound in that solution.

Percentage ionization is the extent of dissociated form of an acid/base in a solution.

Fundamentals

For a solution, pH is equal to the negative of log of concentration of ions in a solution.

pH = -log[H]

For acid HA,

HA—
>H* +A

Acid dissociation constant is calculated as follows:

K -[H] A-7
(HA)

Here, is concentration of hydrogen ion, is concentration of anion and is concentration of undissociated acid in solution.

Percentage ionization is calculated as follows:

001*
[HA]

Here, is concentration of hydrogen ion and is concentration of undissociated acid in solution.

For an acid,

PK, =-log(K.)

Substitute, thus,

-log(K)=4.679

Or,

K
= 10-4.679
= 2.09x10-5

Calculate the concentration of ions from pH as follows:

pH = -log[H]

Substitute, 4.876 for pH thus,

4.876=-log[H]

Or,

[H*] = 10-1876
= 1.33x10-M

The dissociation reaction of acid is as follows;

HA-
> H+ +A

Concentration of ions is1.33x10-M
. ICE table for this reaction is as follows:

HA
H
ICE
Initial concentration
Change in conc.
C-1.33x10-M
1.33x10- M
| 1.33x10- M

Now,

K -[H] A-7
(HA)

Substitute, 1.33x10-M
for , 1.33x10-M
for , C-1.33x10-M
for and 2.09x10-5
for

(1.33x10-9 M)(1.33x10 M)
(c-1.33x10- M)

Rearrange,

(2.09x10^^)(c)-(2.09x10^)(1.33x10-9 M)=1.77x10-ho
(2.09x10-^)(c)-2.78x10-50 =1.77x10-10
(2.09x10-)(c)=4.55x10-10

Or,

c=2.177x10-5 M

Calculate the percentage ionization as follows:

[14]*100
_[H*]x100
с

Substitute,1.33x10-M
for and 2.77x10-5 M
for thus,

1. - 1.33x10-5 M
x100
2.77x10 M
= 61%

Calculate unionized percent of acid(in)
as follows:

Ul. = 100-1.

Substitute, for thus,

UI, =(100-61)%
= 39%

Ans:

The unionized percent of acid is 39 %